OpenStudy (anonymous):

The perimeter of a rectangle is 34 cm. Given that the diagonal is of length 13 cm. Find the dimensions of the rectangle.

OpenStudy (anonymous):

I should be able to help you, despite not being a tutor here. You must be in precalculus?

OpenStudy (anonymous):

Yes I am currently in precalculus.

OpenStudy (anonymous):

Whoa, you're back

OpenStudy (anonymous):

Sweet, okay. First off, you can split the rectangle up into four triangles, right?

OpenStudy (anonymous):

Isn't it more like two? So then it becomes 2 right angle triangles where I can then use the Pythagorean Theorem?

OpenStudy (anonymous):

I think it's easier to use two diagonal lines, because it makes isosceles triangles

OpenStudy (anonymous):

No, I think you're right. NVM <_<;;

OpenStudy (anonymous):

You want to use the Pythagoriean theorem. Also, write out the formula for perimeter. 2l + 2a = 34

OpenStudy (anonymous):

a^2+l^2=13^2

OpenStudy (anonymous):

The trick is to use a substituion. pick either l or a , and solve for it. In this case, I'll pick L

OpenStudy (anonymous):

2l = 34 - 2a (I'm going to reduce this) l = 17 - a

OpenStudy (anonymous):

a^2+(17-a)^2=13^2

OpenStudy (anonymous):

which simplifies to a+(17-a)=13

OpenStudy (anonymous):

Yup. I don't think that's the proper way to substitute in this case. Solve for l on a^2+l^2 = 13^2

OpenStudy (anonymous):

In other words, get l, without a square, by itself

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

not really :(

OpenStudy (anonymous):

Don't worry ^^ I'm here to help

OpenStudy (anonymous):

okay so l^2=(13-a)^2

OpenStudy (anonymous):

l=(13-a)

OpenStudy (anonymous):

Okay. First off, let's get L^2 by itself A^2 + L^2 = 13^2 This changes to L^2 = 169 - A^2

OpenStudy (anonymous):

If you'll recall, L = 17 - a

OpenStudy (anonymous):

So, (17-a)^2 = 169 - A^2

OpenStudy (anonymous):

Do you remember how to foil?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Let's foil (17-a)^2

OpenStudy (anonymous):

289+a^2

OpenStudy (anonymous):

Nope. Remember you have to multiply both the outter and inner terms

OpenStudy (anonymous):

17 * 17, then 17 * -a -a * 17, then -a * -a

OpenStudy (anonymous):

Hmm, I think I need a whiteboard O_O

OpenStudy (anonymous):

i get a=3.52941...

OpenStudy (anonymous):

*does some math (I didn't get that far yet)

OpenStudy (anonymous):

Hmm, that's not what I got. Let's keep going.

OpenStudy (anonymous):

289 - 34a - a^2 = 169 - L^2

OpenStudy (anonymous):

(17-a)(17-a)=-a^2-34a+289

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

We drop that into our equation, and then we need to clean up.

OpenStudy (anonymous):

Bring all of the terms to one side.

OpenStudy (anonymous):

Did I lose you? <_< (Sorry I'm a little slow, I'm tired O_O)

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Okay. (I'm trying not to do all the work for you.) Please bring all of the terms to one side of the equation, and post them. (We're close to being done)

OpenStudy (anonymous):

-a^2-34a+120=-l^2

OpenStudy (anonymous):

(erm, sorry about the -l^2. That's -a^2) (I solved for l)

OpenStudy (anonymous):

so -34a+120=0

OpenStudy (anonymous):

Also, remember that when you square a negative, it becomes positive.

OpenStudy (anonymous):

-34a=-120

OpenStudy (anonymous):

a=3.5294...

OpenStudy (anonymous):

So, it was a^2 - 34a +120 = -a^2 That becomes 2a^2 - 34a + 120

OpenStudy (anonymous):

You forgot that when you multiply a negative by a negative, it becomes positive. (Back when you foiled)

OpenStudy (anonymous):

yeah kinda tired >.>

OpenStudy (anonymous):

It's fine. It's usually a stupid mistake that kills me in math.

OpenStudy (anonymous):

The next step is to reduce. Always reduce whenever you can, because it makes it math easier. (I don't know about you, but I'm lazy when it comes to math:P) So, 2a^2 - 34a + 120 becomes a^2 - 17a + 120

OpenStudy (anonymous):

All you need to do is factor, and you'll get your answer. s. It's plural. You can use the perimeter formula thingie to double check your answer.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

its decomposition...

OpenStudy (anonymous):

Tell what you get ^^

OpenStudy (anonymous):

a=5, 12

OpenStudy (anonymous):

(2a-24)(a-5)

OpenStudy (anonymous):

yeah just too tired... making too many mistakes

OpenStudy (anonymous):

It's fine! You're pro now! Those answers are correct. ^^ 144 + 25 = 169 If you want to impress (cough) your math teacher, you can write the answer as if a = 5, then l = 12, or if a = 12, then l = 5

OpenStudy (anonymous):

The reason either answer works, is because we don't know which way the rectangle is orientated. Oh, and on a funny note, I tried solving the problem with trig. (I'm in Trig right now. That's the math class after precalc) It completely... Did not work. XD Yeesh, I'm tired too. I made some mistakes.

OpenStudy (anonymous):

Anyway, take it easy. God bless with the rest of your homework ^^

OpenStudy (anonymous):

yeah you too :)

OpenStudy (anonymous):

Maybe if I'm lucky, someone will help with my problem. XD